# Cosmic Collider Gravitational Waves sourced by Right-handed Neutrino production from Bubbles: Testing Seesaw, Leptogenesis and Dark Matter

### v1

@@ IV.3.1 Type-I Seesaw
where $M_{N}$ is the mass of the heavy seesaw state and $v$ is the electroweak scale vacuum expectation value (vev). For this work we start with a conventional Type I SeesawMinkowski:1977sc; gellmann2013complexspinorsunifiedtheories; Glashow:1979nm,with three right handed neutrinos NN
@@
To go to a basis where the charged lepton mass matrix is diagonal, we adopt the Casas-Ibarra parametrization Casas_2001 and write the lepton interaction matrix as
@@
Here $\epsilon_{max}$ is related to the Davidson-Ibarra bound Davidson_2002 and is given by
@@
For the hierarchical SM neutrino masses, $m_{3}-m_{1}\sim 0.05$ eV 10.1093/ptep/ptac097. By plugging in this maximum CP violation in the expression for baryon asymmetry (Eq. 65), one gets a lower bound $M_{1}\gtrsim 10^{9}$ GeV. On the other hand, $\epsilon_{max}$ can atmost be 1, which sets an upper bound $M_{1}\lesssim 10^{15}$ GeV. In presence of this CP violation, decays of the RHNs will produce the lepton asymmetry which then would get converted into the baryon asymmetry via Electroweak Sphalerons.
@@ IV.3.2 Non-Thermal production of RHNs and BAU
where $n_{B}$, $n_{\bar{B}}$ and s are the number densities of baryons, antibaryons and entropy at present time. A convenient and simple parameterization of the baryon asymmetry produced in our leptogenesis scenario is cataldi2024leptogenesisbubblecollisions
@@
Here $\mathcal{Y}_{N}$ is the RHN yield produced from bubble collisions, $c_{sph}$ = 28/79 is the sphaleron conversion factor PhysRevD.42.3344, $\epsilon_{CP}$ is the CP violation coming from the decays of RHN to SM given by Eq. (62) and $\kappa_{wash}$ is efficiency factor parameterizing the washout effects from the various processes during leptogenesis. For non-thermal production, other washout effects are not present due to lack an interacting thermal plasma which otherwise would be present in any thermal leptogenesis scenario. For our purposes we work in the limit where wash out processes are negligible, $\kappa_{wash}=1$.
@@ IV.4 Formation of Baryon Asymmetry and its GW probe
The green shaded regions on the right-hand side indicate where the scalar vev becomes comparable to the RHN mass, rendering the baryon-yield calculation unreliable. In this regime the RHNs obtain their mass predominantly from the phase transition, and, as noted earlier, the particle-production formalism break down shakya2025aspectsparticleproductionbubble.
@@
We show in Fig. 11 the detection thresholds for the GW signals associated with the leptogenesis scenario described above. On the same plane, assuming maximal CP violation, we show the region of parameter space that yields sufficient baryon asymmetry for RHN masses in the allowed range $M_{1}\in[10^{9},10^{15}]$ GeV. Although part of the parameter space is already excluded by LVK searches (orange and gray hatched regions), future detectors such as BBO and ET retain significant discovery potential. For example, for a very heavy RHN mass $M_{1}=10^{15}$ GeV and a fast phase transition, $\beta/H\sim 10^{3}$, one expects a bubble-collision signal within the ET sensitivity band together with a RHN-production GW signal in the BBO frequency range at temperatures $T_{*}\sim 10^{7}$ GeV. As the RHN-production contribution exhibits a characteristic peaked spectrum, with a low-frequency rise (approximately $f^{1}$), this again enables one to conclusively probe the RHN-production mechanism and leptogenesis with a inter-detector probing.
@@ V AsDM via RHN decays produced from bubbles
Besides these, current observations show that the DM and Baryon relics are very close today, with $\Omega_{DM}/\Omega_{B}\sim 5$ which is roughly within the same order of magnitude. [^12] Note, as compared to this, for the Dark Energy relic ($\Omega_{\Lambda}$) observed today one has $\Omega_{\Lambda}/\Omega_{b}\gtrsim\mathcal{O}(10)$. This suggests the possibility of a common origin for both relics. One class of scenarios that naturally relates the baryon and dark matter abundances is Asymmetric Dark Matter (AsDM) Nussinov:1985xr; Kaplan:1991ah. In such models, the dark matter particle possesses a distinct antiparticle, and comparable asymmetries in the visible and dark sectors are generated by similar microphysics, for instance through the decay of same heavy particle.
Building on this idea, several works have explored asymmetric dark matter arising from RHN decays during leptogenesis : for $n_{DM}\sim n_{b}$ scenarios, where an asymmetry is first produced in one sector and is then transferred to the other sector at later times via some portal interactions Cosme_2005; Gu_2009 and for a two sector leptogenesis Falkowski:2011xh where asymmetry is generated in both visible and dark sectors simultaneously.
In what follows, we consider a two-sector leptogenesis framework similar to Ref. Falkowski:2011xh. However, unlike the conventional setup, the initial RHN abundance in our scenario is generated non-thermally through production from runaway bubble collisions during a first-order phase transition in the early universe (see Fig. 1). As we show below, this leads to a testable realization of co-genesis, since the associated RHN-production mechanism imprints a distinctive GW signature that can be probed in upcoming experiments.
@@ V.1 Leptogenesis and Asymmetric Dark Matter
In order to generate an asymmetry in leptons and in DM, there must be CP-violation in the decays of $N$. We take the hierarchal approximation Falkowski:2011xh, $M_{1}\ll M_{2}$ and consider only the asymmetry generated by the decays of $N_{1}$ :
@@
where the total decay width for the RHN is $\Gamma_{tot}=\Gamma_{1}+\Gamma_{\chi}$, with $\Gamma_{1}$ defined in Eq. (58) and $\Gamma_{\chi}$ given as Datta:2025vyu
@@
Hence the maximum CP violation and the Davidson-Ibarra bound gets slightly modified Falkowski:2011xh
@@
Then the produced baryon asymmetry and DM yield can be parametrized asDatta:2025vyu
@@ V.2 AsDM formation
The DM yield can be related to the DM mass $m_{DM}$ as below Datta:2025vyu
@@
Assuming a hierarchy between the RHN masses $M_{1}:M_{2}:M_{3}=1:10:100$, we get $\epsilon_{\chi}\lesssim 10^{-2}$ Falkowski:2011xh. We also assume maximal CP violation for $\epsilon_{L}$ given in Eq. (70). This leaves us with a total of three Lagrangian parameters : $M_{1},\ y_{\phi},\ y_{\chi}$ and three P.T. parameters $\alpha,\ \beta/H,\ T_{*}$ that affects the Baryon and DM relic. We trade $y_{\chi}$ for the Branching ratio $\rm{Br}_{\chi}$ and look for the parameter space for AsDM which satisfy the leptogenesis condition as well as active neutrino masses.
@@ VI Viable UV-complete Particle Theory: Multi-Majoron Model
From the observations of various neutrino oscillation experiments, the mass square splittings of the left-handed neutrino mass have been measured as $\Delta m_{21}^{2}\simeq 7.5\times 10^{-5}\mathrm{eV}^{2}$ and $\Delta m_{3\ell}^{2}\simeq 2.5\times 10^{-3}\mathrm{eV}^{2}$ Esteban:2024eli are known. Beside this,the cosmological bound for the sum of left-handed neutrino masses is approaching $0.1$ eV Palanque-Delabrouille:2019iyz; eBOSS:2020yzd. both these indicate the left-handed neutrino mass spectrum is likely to be hierarchical.
The right-handed neutrinos involved in the type I seesaw mechanism are also expected to have a strong hierarchical mass spectrum in several extensions in general, and also motivated from the explaining the puzzle of hierarchial fermion masses in the SM Froggatt:1978nt; Barr:1979xt; Barbieri:1983uq; Balakrishna:1987qd; Babu:1988fn; Babu:2004th; Babu:2009fd, see Refs. Babu:1999me; Yoshioka:2000tve; Chao:2012re; Altmannshofer:2014qha; Higuchi:2014cua; Huitu:2017ukq; Okada:2019fgm; Hernandez:2021iss; Bonilla:2023wok; FernandezNavarro:2024hnv; Arbelaez:2024rbm; Jana:2024icm for attempts to bring the seesaw explanation of neutrino mass and fermion mass hierarchy under one umbrella  [^13] Models with sequential dominance King:2013eh; Fu:2023nrn or vanilla leptogenesis Buchmuller:2004nz; Davidson:2008bu also predict such hierarchical RHN masses..
In this section, as an illustrative example only, we consider a minimal extension of the SM explaining the hierarchical right-handed neutrino masses is what is known as multi-Majoron models Bamert:1994hb, where different right-handed neutrinos $N_{i}$ couple to a complex scalar Majoron fields $\phi_{i}$ through Yukawa interaction which we describe in detail below. The model we is based on a global ${\rm U}(1)_{B-L}\times\rm{U}(1)_{N}$ of the SM gauge group, $SU(3)_{c}\times SU(2)_{L}\times U(1)_{Y}$ Fu:2025dlp; DiBari:2023mwu.
@@
Three generations of RHNs $N_{i}$ ($i=1,2,3$) are introduced along with two additional complex scalar fields $\Phi_{1},\Phi_{2}$, charged under the global $U(1)_{B-L}$ and $U(1)_{N}$ respectively. These scalars acquiring vev spontaneously breaks the $U(1)_{B-L}\times U(1)_{N}$ symmetry, which contribute to the masses of the RHNs. The particle content of the model is listed in Table 1. The additional Yukawa terms involving the RHNs are given by  DiBari:2023mwu:
@@
where the first term gives the Dirac neutrino mass after electroweak symmetry breaking, while the next terms generate the RHN Majorana masses $M_{i}$. We write the complex scalars as $\Phi_{1}=\phi_{1}e^{i\theta_{1}}/\sqrt{2}$ and $\Phi_{2}=\phi_{2}e^{i\theta_{2}}/\sqrt{2}$ and take the vev along the real axis, $\langle\Phi_{1}\rangle=v_{1}/\sqrt{2},\ \langle\Phi_{2}\rangle=v_{2}/\sqrt{2}$. We further assume the hierarchy $v_{1}\gg v_{2}$ such that the Majorana masses satisfy $M_{3}\gg M_{1}\approx M_{2}\approx M$. Note this masses enter in the seesaw expression in Eq. (52) and give the SM neutrino masses : $m_{\nu_{1,2}}\sim v_{h}^{2}/M,\ m_{\nu_{3}}\sim v_{h}^{2}/M_{3}$, $v_{h}$ being the Higgs vev. Thus Multi-Majoron model seesaw generates a hierarchy between the SM neutrino masses as well which is different than standard single Majoron. Since $M\sim y_{1}v_{2}/2,\ M_{3}\sim y_{3}v_{1}/2$, the vevs $v_{1},v_{2}$ of the $U(1)_{N}$ and $U(1)_{B-L}$ breaking scalars $\Phi_{2},\Phi_{1}$ play important role in the seesaw mechanism which can further be probed via the novel GW signature.
Along with the usual quadratic and quartic terms, the symmetry $\rm U(1)_{N}\times\rm U(1)_{B-L}$ also allows a quartic mixing term and the tree level potential can be written as
@@
At high temperatures both symmetries are restored. The $U(1)_{B-L}$ symmetry is broken at $T\sim v_{1}$, generating the mass $M_{3}$ for the heaviest RHN. GWs produced during this transition are negligible DiBari:2023mwu. After $\Phi_{1}$ settles at its vev, the $\Phi_{2}$ phase transition occurs. Owing to the scalar mixing term, the zero-temperature effective potential for $\phi_{2}$ develops a cubic term in the mass-diagonal basis [^14] See Eq. (3.8) of Ref. DiBari:2023mwu. : $\{\varphi_{1},\varphi_{2}\}$, leading to a strongly first-order phase transition.
@@
which for our case reduces to DiBari:2023mwu
@@
where $m_{\varphi_{2}}$ is the mass of the scalar $\varphi_{2}$ after $\phi_{1}$ and $\phi_{2}$ both acquires their VEVs : $m_{\varphi_{2}}^{2}=-\mu_{2}^{2}-6Cv_{2}+3\lambda_{2}v_{2}^{2}\ ;$ $a_{F},a_{B}$ are constants arising due to the high-temperature expansion of the fermionic / bosonic thermal function and
@@
Next, from this effective thermal potential we calculate the $\mathcal{O}(3)$ symmetric Euclidean action $S_{3}$ Coleman:1977py,Linde:1981zj
@@
with the boundary conditions $\left(d\varphi_{2}/dr\right)_{r=0}=0$ and $\varphi_{2}\left(r\rightarrow\infty\right)=0$. It can be shown Megevand:2016lpr that at the phase transition temperature $T_{*}$ the Euclidean action $S_{3}$ has to satisfy [^15] It may be quite possible that in certain regions of the parameter space, $\mathcal{O}(4)$ bounce solution dominates over $\mathcal{O}(3)$ and the Euclidean action cannot be taken as $S_{E}=S_{3}/T$. However we expect such regions to be small Ghoshal:2020vud. Such an analysis is beyond the scope of the present paper.
@@
During the $\varphi_{2}$ phase transition, collisions of expanding false-vacuum bubbles can result in production of $\varphi_{1}$ particles. The produced density will be non-thermal as $m_{\varphi_{1}}\sim\sqrt{2\lambda_{1}}v_{1}\gg v_{2}\sim T_{*}$ which can potentially source leptogenesis alongside a characteristic stochastic GW signal. In this work, we realize this mechanism using the scalar-portal setup of Ref. cataldi2024leptogenesisbubblecollisions, unlike the RHN-portal framework employed in Secs. IV and V. In the scalar-portal setup, an initial density of the scalar $\varphi_{1}$ particles is produced from the $\varphi_{2}$ bubble collisions which subsequently decays into the RHNs.
@@
$\kappa_{\phi 2}\approx\frac{3}{64\pi^{5}}\frac{\zeta^{2}}{\lambda}\ln[\dots]\left(1+\mathcal{O}\left(M_{pl}^{-1}\right)\right).$ (84)
@@
Figure 14: Three Benchmark points $BP_{X}$, $BP_{Y}$ and $BP_{Z}$ for different values of $\lambda_{2}$, $\zeta$,$M$ and $v_{2}$ in the $U(1)_{N}\times U(1)_{B-L}$ model. The shaded regions corresponds to SNR >10>10 for various GW experiments. $\alpha=400$ and $\kappa_{\phi 2}\sim\mathcal{O}(1)$ taken for $\phi_{1}$ production.
Next we calculate the phase transition parameters $\alpha,\ \beta/H$ and $T_{*}$ from the potential in Eq. (78) by evaluating the Euclidean action $S_{3}$ and using the relations Eq. (81-83). By performing a random parameter scan of the variables $\{\lambda_{2},v_{2},\zeta,M\}$ we note down three benchmark points $BP_{X}$, $BP_{Y}$ and $BP_{Z}$ in Fig. 14. Corresponding values of different coupling strengths, masses and P.T. parameters are summarized in Table. 2.
@@
Fig. 14 shows for all three Benchmarks, the GWs produced from the corresponding phase transition can be probed in detectors BBO and ET with high SNR. This will be possible for GWs coming from the novel particle production mechanism alongside with the well established bubble collision GWs  [^17] Besides there can also be GW arising due to global cosmic strings from global $U(1)_{\rm B-L}$ but those are hardly in the detectable range for the choice of $v_{1}$ and moreover the estimates for GW from global cosmic strings have huge uncertainty and may vary orders of magnitude, so we do not discuss them in this analysis, see Refs. Fu:2023nrn; Fu:2025dlp for some detail.. As was shown in Fig. 2, the RHN production sources GW signals which have a unique characteristic tail of $\Omega_{GW}h^{2}\sim f^{1}$ for lower frequencies and hence can be separated from the GW signals coming from bubble collisions which has the form approximate $\Omega_{GW}h^{2}\sim f^{2.3}$.
@@
For each benchmark point we also calculate the corresponding BAU and DM relics for the three scenarios discussed in sec IV&V. The yield of the produced $N_{3}$ number density in the scalar portal setup is given as (see Appendix-A and Ref. cataldi2024leptogenesisbubblecollisions )

### v2

@@ IV.3.1 Type-I Seesaw
where $M_{N}$ is the mass of the heavy seesaw state and $v$ is the electroweak scale vacuum expectation value (vev). For this work we start with a conventional Type I SeesawMinkowski (1977); Gell-Mann et al. (2013); Glashow (1980),with three right handed neutrinos NN
@@
To go to a basis where the charged lepton mass matrix is diagonal, we adopt the Casas-Ibarra parametrization Casas and Ibarra (2001) and write the lepton interaction matrix as
@@
Here $\epsilon_{max}$ is related to the Davidson-Ibarra bound Davidson and Ibarra (2002) and is given by
@@
For the hierarchical SM neutrino masses, $m_{3}-m_{1}\sim 0.05$ eV Group et al. (2022). By plugging in this maximum CP violation in the expression for baryon asymmetry (Eq. 65), one gets a lower bound $M_{1}\gtrsim 10^{9}$ GeV. On the other hand, $\epsilon_{max}$ can atmost be 1, which sets an upper bound $M_{1}\lesssim 10^{15}$ GeV. In presence of this CP violation, decays of the RHNs will produce the lepton asymmetry which then would get converted into the baryon asymmetry via Electroweak Sphalerons.
@@ IV.3.2 Non-Thermal production of RHNs and BAU
where $n_{B}$, $n_{\bar{B}}$ and s are the number densities of baryons, antibaryons and entropy at present time. A convenient and simple parameterization of the baryon asymmetry produced in our leptogenesis scenario is Cataldi and Shakya (2024)
@@
Here $\mathcal{Y}_{N}$ is the RHN yield produced from bubble collisions, $c_{sph}$ = 28/79 is the sphaleron conversion factor Harvey and Turner (1990), $\epsilon_{CP}$ is the CP violation coming from the decays of RHN to SM given by Eq. (62) and $\kappa_{wash}$ is efficiency factor parameterizing the washout effects from the various processes during leptogenesis. For non-thermal production, other washout effects are not present due to lack an interacting thermal plasma which otherwise would be present in any thermal leptogenesis scenario. For our purposes we work in the limit where wash out processes are negligible, $\kappa_{wash}=1$.
@@ IV.4 Formation of Baryon Asymmetry and its GW probe
The green shaded regions on the right-hand side indicate where the scalar vev becomes comparable to the RHN mass, rendering the baryon-yield calculation unreliable. In this regime the RHNs obtain their mass predominantly from the phase transition, and, as noted earlier, the particle-production formalism break down Shakya (2025c).
@@
We show in Fig. 11 the detection thresholds for the GW signals associated with the leptogenesis scenario described above. On the same plane, assuming maximal CP violation, we show the region of parameter space that yields sufficient baryon asymmetry for RHN masses in the allowed range $M_{1}\in[10^{9},10^{15}]$ GeV. Although part of the parameter space is already excluded by LVK searches (orange and gray hatched regions), future detectors such as BBO, ET and future LVK observations retain significant discovery potential. For example, for a very heavy RHN mass $M_{1}=10^{15}$ GeV and a fast phase transition, $\beta/H\sim 10^{3}$, one expects a bubble-collision signal within the ET and future LVK sensitivity bands together with a RHN-production GW signal in the BBO frequency range at temperatures $T_{*}\sim 10^{7}$ GeV. As the RHN-production contribution exhibits a characteristic peaked spectrum, with a low-frequency rise (approximately $f^{1}$), this again enables one to conclusively probe the RHN-production mechanism and leptogenesis with a inter-detector probing.
@@ V AsDM via RHN decays produced from bubbles
Besides these, current observations show that the DM and Baryon relics are very close today, with $\Omega_{DM}/\Omega_{B}\sim 5$ which is roughly within the same order of magnitude. [^12] Note, as compared to this, for the Dark Energy relic ($\Omega_{\Lambda}$) observed today one has $\Omega_{\Lambda}/\Omega_{b}\gtrsim\mathcal{O}(10)$. This suggests the possibility of a common origin for both relics. One class of scenarios that naturally relates the baryon and dark matter abundances is Asymmetric Dark Matter (AsDM) Nussinov (1985); Kaplan (1992). In such models, the dark matter particle possesses a distinct antiparticle, and comparable asymmetries in the visible and dark sectors are generated by similar microphysics, for instance through the decay of same heavy particle.
Building on this idea, several works have explored asymmetric dark matter arising from RHN decays during leptogenesis : for $n_{DM}\sim n_{b}$ scenarios, where an asymmetry is first produced in one sector and is then transferred to the other sector at later times via some portal interactions Cosme et al. (2005); Gu et al. (2009) and for a two sector leptogenesis Falkowski et al. (2011b) where asymmetry is generated in both visible and dark sectors simultaneously.
In what follows, we consider a two-sector leptogenesis framework similar to Ref. Falkowski et al. (2011b). However, unlike the conventional setup, the initial RHN abundance in our scenario is generated non-thermally through production from runaway bubble collisions during a first-order phase transition in the early universe (see Fig. 1). As we show below, this leads to a testable realization of co-genesis, since the associated RHN-production mechanism imprints a distinctive GW signature that can be probed in upcoming experiments.
@@ V.1 Leptogenesis and Asymmetric Dark Matter
In order to generate an asymmetry in leptons and in DM, there must be CP-violation in the decays of $N$. We take the hierarchal approximation Falkowski et al. (2011b), $M_{1}\ll M_{2}$ and consider only the asymmetry generated by the decays of $N_{1}$ :
@@
where the total decay width for the RHN is $\Gamma_{tot}=\Gamma_{1}+\Gamma_{\chi}$, with $\Gamma_{1}$ defined in Eq. (58) and $\Gamma_{\chi}$ given as Datta et al. (2025)
@@
Hence the maximum CP violation and the Davidson-Ibarra bound gets slightly modified Falkowski et al. (2011b)
@@
Then the produced baryon asymmetry and DM yield can be parametrized asDatta et al. (2025)
@@ V.2 AsDM formation
The DM yield can be related to the DM mass $m_{DM}$ as below Datta et al. (2025)
@@
Assuming a hierarchy between the RHN masses $M_{1}:M_{2}:M_{3}=1:10:100$, we get $\epsilon_{\chi}\lesssim 10^{-2}$ Falkowski et al. (2011b). We also assume maximal CP violation for $\epsilon_{L}$ given in Eq. (70). This leaves us with a total of three Lagrangian parameters : $M_{1},\ y_{\phi},\ y_{\chi}$ and three P.T. parameters $\alpha,\ \beta/H,\ T_{*}$ that affects the Baryon and DM relic. We trade $y_{\chi}$ for the Branching ratio $\rm{Br}_{\chi}$ and look for the parameter space for AsDM which satisfy the leptogenesis condition as well as active neutrino masses.
@@ VI Viable UV-complete Particle Theory: Multi-Majoron Model
From the observations of various neutrino oscillation experiments, the mass square splittings of the left-handed neutrino masses have been measured as $\Delta m_{21}^{2}\simeq 7.5\times 10^{-5}\mathrm{eV}^{2}$ and $\Delta m_{3\ell}^{2}\simeq 2.5\times 10^{-3}\mathrm{eV}^{2}$ Esteban et al. (2024). Beside this, the cosmological bound for the sum of left-handed neutrino masses is approaching $0.1$ eV Palanque-Delabrouille et al. (2020); Alam and others (2021). Both these indicate the left-handed neutrino mass spectrum is likely to be hierarchical.
The right-handed neutrinos involved in the type I seesaw mechanism are also expected to have a strong hierarchical mass spectrum in several extensions in general, and also motivated from the explaining the puzzle of hierarchial fermion masses in the SM Froggatt and Nielsen (1979); Barr (1980); Barbieri et al. (1983); Balakrishna (1988); Babu and He (1989); Babu and Enkhbat (2005); Babu (2010), see Refs. Babu and Nandi (2000); Yoshioka (2000); Chao (2016); Altmannshofer et al. (2014); Higuchi and Yamamoto (2014); Huitu and Koivunen (2018); Okada et al. (2019); Hernández et al. (2021); Bonilla et al. (2023); Fernández Navarro et al. (2024); Arbeláez et al. (2024); Jana et al. (2025) for attempts to bring the seesaw explanation of neutrino mass and fermion mass hierarchy under one umbrella  [^13] Models with sequential dominance King and Luhn (2013); Fu et al. (2023) or vanilla leptogenesis Buchmuller et al. (2005); Davidson et al. (2008) also predict such hierarchical RHN masses..
In this section, as an illustrative example only, we consider a minimal extension of the SM explaining the hierarchical right-handed neutrino masses in what is known as multi-Majoron models Bamert et al. (1995), where different right-handed neutrinos $N_{i}$ couple to complex scalar Majoron fields $\phi_{i}$ through Yukawa interaction which we describe in detail below. The model we consider is based on a global ${\rm U}(1)_{B-L}\times\rm{U}(1)_{N}$ extension of the SM gauge group, $SU(3)_{c}\times SU(2)_{L}\times U(1)_{Y}$ Fu et al. (2025); Di Bari et al. (2024).
@@
Three generations of RHNs $N_{i}$ ($i=1,2,3$) are introduced along with two additional complex scalar fields $\Phi_{1},\Phi_{2}$, charged under the global $U(1)_{B-L}$ and $U(1)_{N}$ respectively. These scalars acquiring vev spontaneously breaks the $U(1)_{B-L}\times U(1)_{N}$ symmetry, which contribute to the masses of the RHNs. The particle content of the model is listed in Table 1. The additional Yukawa terms involving the RHNs are given by  Di Bari et al. (2024):
@@
where the first term gives the Dirac neutrino mass after electroweak symmetry breaking, while the next terms generate the RHN Majorana masses $M_{i}$. We write the complex scalars as $\Phi_{1}=\phi_{1}e^{i\theta_{1}}/\sqrt{2}$ and $\Phi_{2}=\phi_{2}e^{i\theta_{2}}/\sqrt{2}$ and take the vev along the real axis, $\langle\Phi_{1}\rangle=v_{1}/\sqrt{2},\ \langle\Phi_{2}\rangle=v_{2}/\sqrt{2}$. We further assume the hierarchy $v_{1}\gg v_{2}$ such that the Majorana masses satisfy $M_{3}\gg M_{1}\approx M_{2}\approx M$. Note this masses enter in the seesaw expression in Eq. (52) and give the SM neutrino masses : $m_{\nu_{1,2}}\sim v_{h}^{2}/M,\ m_{\nu_{3}}\sim v_{h}^{2}/M_{3}$, $v_{h}$ being the Higgs vev. Thus Multi-Majoron model seesaw generates a hierarchy between the SM neutrino masses as well, which is different than standard single Majoron. Since $M\sim y_{1}v_{2}/2,\ M_{3}\sim y_{3}v_{1}/2$, the vevs $v_{1},v_{2}$ of the $U(1)_{N}$ and $U(1)_{B-L}$ breaking scalars $\Phi_{2},\Phi_{1}$ play important role in the seesaw mechanism which can further be probed via the novel GW signature.
Along with the usual quadratic and quartic terms, the symmetry $\rm U(1)_{N}\times\rm U(1)_{B-L}$ also allows a quartic mixing term in the tree level potential and can be written as
@@
At high temperatures both symmetries are restored. The $U(1)_{B-L}$ symmetry is broken at $T\sim v_{1}$, generating the mass $M_{3}$ for the heaviest RHN $N_{3}$. GWs produced during this transition are negligible Di Bari et al. (2024). After $\Phi_{1}$ settles at its vev, the $\Phi_{2}$ phase transition occurs. Owing to the scalar mixing term, the zero-temperature effective potential for $\phi_{2}$ develops a cubic term in the mass-diagonal basis [^14] See Eq. (3.8) of Ref. Di Bari et al. (2024). : $\{\varphi_{1},\varphi_{2}\}$, leading to a strongly first-order phase transition.
@@
which for our case reduces to Di Bari et al. (2024)
@@
where $m_{\varphi_{2}}$ is the mass of the scalar $\varphi_{2}$ after $\phi_{1}$ and $\phi_{2}$ both acquires their VEVs
$m_{\varphi_{2}}^{2}=-\mu_{2}^{2}-6Cv_{2}+3\lambda_{2}v_{2}^{2}\ ;$
$a_{F},a_{B}$ are constants arising due to the high-temperature expansion of the fermionic / bosonic thermal function and
@@
Next, from this effective thermal potential we calculate the Euclidean action $S_{E}$. Note for temperature dependent potentials the least action solution has $\mathcal{O}(3)$ symmetry and $S_{E}=S_{3}/T$ where $S_{3}$ Coleman (1977),Linde (1983) is the $\mathcal{O}(3)$ symmetric Euclidean action
@@
with the boundary conditions $\left(d\varphi_{2}/dr\right)_{r=0}=0$ and $\varphi_{2}\left(r\rightarrow\infty\right)=0$. It can be shown Megevand and Ramirez (2017) that at the phase transition temperature $T_{*}$ the Euclidean action $S_{3}$ has to satisfy
@@
Before we move to calculate the phase transition parameters using the above equations, we note that for certain regions of parameter space of our potential in Eq. (76), the decay rate of the false vacuum might be dominated by the $\mathcal{O}(4)$ symmetric Euclidean action $S_{4}$ instead of the $\mathcal{O}(3)$ symmetric action $S_{3}$ and in general one has to consider Ghoshal and Salvio (2020):
$\Gamma=\max\left\{T^{4}\left(\frac{S_{3}}{2\pi T}\right)^{3/2}\exp(-S_{3}/T),\;\frac{1}{R_{4}^{4}}\left(\frac{S_{4}}{2\pi}\right)^{2}\exp(-S_{4})\right\},$ (84)
where $\Gamma$ is the decay rate of the false vacuum and $R_{4}$ is the radius of bubbles at nucleation for the $\mathcal{O}(4)$ bounce.
During the $\varphi_{2}$ phase transition, collisions of expanding false-vacuum bubbles can result in production of $\varphi_{1}$ particles. The produced density will be non-thermal as $m_{\varphi_{1}}\sim\sqrt{2\lambda_{1}}v_{1}\gg v_{2}\sim T_{*}$ which can potentially source leptogenesis alongside a characteristic stochastic GW signal. In this work, we realize this mechanism using the scalar-portal setup of Ref. Cataldi and Shakya (2024), unlike the RHN-portal framework employed in Secs. IV and V. In the scalar-portal setup, an initial density of the scalar $\varphi_{1}$ particles is produced from the $\varphi_{2}$ bubble collisions which subsequently decays into the RHNs.
@@
$\kappa_{\varphi 1}\approx\frac{3}{64\pi^{5}}\frac{\zeta^{2}}{\lambda}\ln[\dots]\left(1+\mathcal{O}\left(M_{pl}^{-1}\right)\right).$ (85)
@@
Figure 14: Three Benchmark points $BP_{X}$, $BP_{Y}$ and $BP_{Z}$ for different values of $\lambda_{2}$, $\zeta$, $M$ and $v_{2}$. The shaded regions corresponds to SNR >10>10 for various GW experiments. The hatched region is ruled out by current LVK observations. $\alpha=400$ and $\kappa_{\varphi 1}\sim\mathcal{O}(1)$ taken for $\varphi_{1}$ production.
Next we calculate the phase transition parameters $\alpha,\ \beta/H$ and $T_{*}$ from the potential in Eq. (78) by evaluating the Euclidean action $S_{3}$ and using the relations Eq. (81-83). By performing a random parameter scan of the variables $\{\lambda_{2},v_{2},\zeta,M\}$ we note down three benchmark points [^16] We have checked explicitly for the benchmark points $S_{4}\gg S_{3}/T$. Using Eq. (84) we then take $S_{E}=S_{3}/T$ for bounce calculation. $BP_{X}$, $BP_{Y}$ and $BP_{Z}$ in Fig. 14. Corresponding values of different coupling strengths, masses and P.T. parameters are summarized in Table. 2.
@@
Fig. 14 shows for all three Benchmarks, the GWs produced from the corresponding phase transition can be probed in detectors BBO and ET with high SNR. This will be possible for GWs coming from the novel particle production mechanism alongside with the well established bubble collision GWs  [^17] Besides there can also be GW arising due to global cosmic strings from global $U(1)_{\rm B-L}$ but those are hardly in the detectable range for the choice of $v_{1}$ and moreover the estimates for GW from global cosmic strings have huge uncertainty and may vary orders of magnitude, so we do not discuss them in this analysis, see Refs. Fu et al. (2023, 2025) for some detail.. As was shown in Fig. 2, the RHN production sources GW signals which have a unique characteristic tail of $\Omega_{GW}h^{2}\sim f^{1}$ for lower frequencies and hence can be separated from the GW signals coming from bubble collisions which has the form approximate $\Omega_{GW}h^{2}\sim f^{2.3}$.
@@
For each benchmark point we also calculate the corresponding BAU and DM relics for the three scenarios discussed in sec IV&V. The yield of the produced $N_{3}$ number density in the scalar portal setup is given as (see Appendix-A and Ref. Cataldi and Shakya (2024) )